OpenStudy (anonymous):

a (n)=4n+4 .WRITE FIRST 4 TERMS OF THE SEQUECE WHOSE GENERAL TERM IS GIVIN

OpenStudy (anonymous):

find: a(0) a(1) a(2) a(3)

OpenStudy (anonymous):

\[a(0)=4\times 0+4=4\] \[a(1)=4\times 1+4=8\] others are similar